# Robotics meets Fluid Dynamics: A Characterization of the Induced Airflow around a Quadrotor

### v1

@@ Abstract
The widespread adoption of quadrotors for diverse applications, from agriculture to public safety, necessitates an understanding of the aerodynamic disturbances they create. This paper introduces a computationally lightweight model for estimating the time-averaged magnitude of the induced flow below quadrotors in hover. Unlike related approaches that rely on expensive computational fluid dynamics (CFD) simulations or time-consuming empirical measurements, our method leverages classical theory from turbulent flows. By analyzing over 9 hours of flight data from drones of varying sizes within a large motion capture system, we show that the combined flow from all propellers of the drone is well-approximated by a turbulent jet. Through the use of a novel normalization and scaling, we have developed and experimentally validated a unified model that describes the mean velocity field of the induced flow for different drone sizes. The model accurately describes the far-field airflow in a very large volume below the drone which is difficult to simulate in CFD. Our model, which requires only the drone’s mass, propeller size, and drone size for calculations, offers a practical tool for dynamic planning in multi-agent scenarios, ensuring safer operations near humans and optimizing sensor placements.
@@ I Introduction
In recent years, quadrotors have gained popularity for a wide variety of tasks in academia [1, 2, 3, 4] as well as in industry where companies develop drones for filming [5], mapping [6], inspection [7], and public safety [8]. Shared across these diverse applications is the need to understand and characterize the strong downwash generated by the vehicle’s propellers as this enables informed decisions on the allowed proximity of a quadrotor to an object or person, ultimately leading to safer autonomous drones.
@@
In mapping and inspection tasks this helps understanding when aerodynamic interactions with close-by structures such as bridges, powerlines [9], or ships [10] occur. When drones are operated for agricultural purposes, it is critical to know how far the aerodynamic disturbances caused by the drone extend for plant spraying and protection [11]. When deployed for filming and in public spaces, quadrotors are often operated in the vicinity of people and minimizing the presence of intrusive flows in the scene is important. In scenarios where multiple drones are operated together, a flow model can be used to improve planning such that individual vehicles dynamically avoid each other’s downwash and show an improved response to external disturbances [12]. Finally, a better understanding of the flow can be important for sensor and scientific instrumentation placement [13, 14, 15, 16, 17].
This paper presents a computationally lightweight approach to model the time-averaged magnitude of the induced flow around a quadrotor at hover. Put differently, it answers the question: ‘How much wind does a drone generate when hovering or flying slowly in a near-hover state?’
Figure 1: A smoke visualization of the near-field flow at hover where the flow contributions of the individual propellers are separated. Further downstream they merge into one turbulent jet (far-field).
@@
Detailed computation of the induced flow around a drone is a very challenging problem as it requires expensive and time-consuming simulations involving computational fluid dynamics (CFD). To limit the computational demand, CFD simulations typically focus on a very short time-scale and are limited to only simulate the airflow close to the vehicle [18, 19, 20]. This methodology is suitable to assess the performance of the drone, but not suited to analyze the far-field which extends over many meters away from the drone where one needs to rely on turbulence modeling [15, 17]. Additionally, a CFD-based approach is not suitable for a real-time planner to dynamically coordinate a fleet of multiple quadrotors, such that the vehicles minimize interference with each other.
@@
Orthogonal to the simulation approach, fluid dynamics research has a century-old history of heavily relying on empirical measurement. However, experimentally characterizing each drone with thousands of measurements in the entire 3D space around it is a tedious process. Firstly, performing experiments requires a wind-free (indoor) measurement domain much larger than the drone itself to properly identify the aerodynamic signature downstream. Secondly, a proper measurement technique needs to be selected. For example Particle Image Velocimetry (PIV) [21] is well suited, but generally presents additional challenges in large-scale real-world deployments [22, 23].
@@
In this paper, we present a unified, computationally lightweight model to calculate the mean velocity field of the induced flow around a drone at hover. The unified model is inspired by decades of well-established research on turbulent jets [24]. The advance in this work is enabled by joining methodologies from fluid dynamics and robotics research. Recording over 9 h of manual and autonomous drone flight data with three differently sized drones ranging from 572 g to 6.3 kg we perform pointwise flow measurements in a very large motion capture system [25] to acquire thousands of data samples of the mean velocity magnitude. Through the use of appropriate normalization and the introduction of a characteristic length, the model is unified and can be applied to quadrotors of different mass and size.
Our methodology is computationally lightweight as it does not rely on CFD simulations. Instead, it uses available closed-form analytic flow solutions. All calculations only require knowledge of the vehicle’s mass, the vehicle’s dimensions, and the propeller size.
@@ II Related Work
For control and simulation tasks the primary goal of a model is to predict the motion of the vehicle given a certain actuation, that is, predict the forces and torques acting on the vehicle. The most widespread models are quadratic thrust and drag models [26, 27, 28] and more advanced models are based on blade-element-momentum (BEM) theory which calculates the lift and drag a propeller based on classical theory for airfoils [29, 30, 31, 32, 33, 34, 35]. However, these approaches only estimate the aerodynamic forces and torques of the propeller and do not compute the volumetric flow induced by the vehicle.
In contrast to the above models, computational fluid dynamics simulations calculate the three-dimensional flow around the vehicle. Such approaches are typically used to determine and improve the efficiency of the design of a single propeller [36, 37], but have also been applied to simulate entire quadrotors [20, 19], and larger hexacopters [38, 15]. While such simulations achieve results that are highly accurate and manage to capture real-world effects well [39, 40], they require large amounts of computation. Additionally, they do not generalize to other vehicles, requiring a new analysis for every drone.
In contrast to a purely simulation-based approach, a few works perform real-world experiments. For example, [13] performed flow probe measurements on a hexacopter that seem to find overall agreement with later CFD work [15]. In [41] the authors present experimental data to understand the influence of the quadrotor on atmospheric temperature and pressure measurements. In [42] the authors focus on the development of a Schlieren-photography method to qualitatively visualize the flow around drones. In [14] the authors performed outdoor smoke visualization, and in [43] the authors applied PIV to smoke visualizations to estimate the velocity field around a quadrotor model in ground effect. Only a few other studies have applied PIV to drones. For example, in [44] the interaction between propellers under different configurations has been studied, and in [45] the near field flow of a quadrotor model in forward flight is analyzed using wind tunnel experiments. Most recently  [11] performed PIV in a different application and recovered the spreading rate of spray droplets in the drone’s downwash.
Apart from [14, 41], all of these works performed measurements in tethered flight, presenting actual limitations in recovering the true free flight physics. To the authors’ best knowledge, no studies have looked into a comprehensive analysis of the flow around drones in hover while focusing on a scaling analysis, and on closed-form solutions from turbulent flows.
@@ III Theory
This section gives an overview of the relevant notation, overall drone thrust and propeller downwash in hovering flight, and introduces the concept of a free jet. Finally, we introduce several normalizations enabling the analysis to generalize across a wide variety of drones.
Figure 2: Diagram for the drone and the flow coordinate system and geometry. The bodyframe $\mathcal{B}$ of a quadrotor is such that the x-axis faces forward, and the z-axis upwards in thrust direction. The flow coordinate system is oriented such that its longitudinal axis $\mathbf{\hat{s}_{\mathcal{F}}}$ is aligned with flow direction (i.e., points in negative $\mathbf{\hat{z}_{\mathcal{B}}}$ direction). The flow is described in cylindrical coordinates $[s~{}~{}r~{}~{}\theta]^{\top}$ by its longitudinal (axial), radial, and azitmutal velocity components $U$, $V$ and $W$, respectively. The vehicle’s propeller diameter $d$, motor-distance $w$ and tip-to-tip length $l$ are defined as depicted.
@@ III-A Coordinate System and Notation
Figure 2 shows the quadrotor-centered bodyframe coordinate system $\mathcal{B}$, with vector $\mathbf{x}_{\mathcal{B}}=[x~{}~{}y~{}~{}z]^{\top}$. In the bodyframe the $\mathbf{\hat{x}_{\mathcal{B}}}$-axis points forward, the $\mathbf{\hat{y}_{\mathcal{B}}}$-axis points to the left, and the $\mathbf{\hat{z}_{\mathcal{B}}}$-axis points upwards. Since we only consider the quadrotor at hover, the $\mathbf{\hat{z}_{\mathcal{B}}}$-axis is opposing the gravity vector at all times and no rotation due to the roll and pitch of the vehicle need be considered. We also introduce a flow coordinate system $\mathcal{F}$ with $\mathbf{x}_{\mathcal{F}}=[s~{}~{}r~{}~{}\theta]^{\top}$ in cylindrical coordinates along the flow direction $\mathbf{\hat{s}_{\mathcal{F}}}$. The longitudinal coordinate $s$ is aligned along the downwash direction and the radial coordinate $r$ points outwards as shown in Fig. 2. In this coordinate system the velocity vector $\mathbf{U}$ with components $U,V,W$ is defined as shown in Fig. 2.
@@ III-B Propeller-Induced Flow
In a steady-state hover, the multicopter must produce enough thrust to support its weight. This means that—considering a quadrotor—the $N_{r}=4$ propellers must produce a thrust force that is equal to $mg$ where $m$ is the mass and $g$ is the gravitational acceleration.
@@
$U_{\rm{H}}=\sqrt{\frac{T_{\rm{H}}}{2\rho A_{\text{prop}}}}=\sqrt{\frac{\vphantom{T}mg}{2\rho\pi r_{\text{prop}}^{2}N_{r}}}\ ,$ (1)
@@ III-C Turbulent Jet Flow
Here $U_{\text{C}}$ describes the centerline velocity at $r=0$ as function of the distance $s$ below the rotor plane. For a jet with half-width $r_{1/2}$, the rescaled radial position $\xi$ is given by:
$\xi=\frac{r}{r_{1/2}(s)}\ .$ (3)
The jet centerline velocity scales inversely proportional to the distance to the rotor plane. It can be calculated as
@@
A key characteristic of the turbulent jet is that the spreading angle is commonly around 12 deg and is weakly dependent on the flow Reynolds number [24], making the model applicable to a wide range of drone flow and geometry. Furthermore, we remark that we do not need to consider propeller swirl in the analysis as the rotors counter-rotate, and we assume that all other flow components are entrained (drawn/sucked into the flow from the side) downstream.
@@ III-D Normalization
To develop a generalized model of the airflow around different drones, the different phyiscal properties of those drones must be taken into account. We achieve this through introducing a velocity and a spatial normalization based on the theoretical considerations described previously.
@@
From (2) and (4) we observe that the overall scaling of the jet velocity is given by the jet-exit velocity $U_{\text{J}}$. The jet exit velocity is a quantity that describes how fast a turbulent jet exits the nozzle and captures how much momentum the flow carries. While a quadrotor does not have a nozzle, the induced velocity $U_{\rm{H}}$ of eq. (1) relates to the same physical quantity. Denoting a normalized quantity with $\sim$, the normalized velocity $\tilde{U}$ and normalized centerline velocity $\tilde{U}_{C}$ are given by:
$\langle\tilde{U}\rangle(\xi)=\frac{\langle U\rangle(\xi)}{U_{\rm{H}}}\quad\text{and}\quad\tilde{U}_{\text{C}}(s)=\frac{U_{\text{C}}(s)}{U_{\rm{H}}}\;.$ (6)
@@
Similar to the velocity normalization, a spatial normalization parameter is needed to be able to compare drones of different sizes. We propose the tip-to-tip length $l$ as the length-scale parameter as this distance is closest equivalent to the nozzle diameter in a jet flow. Using this, we define the normalized distance to the rotor plane $\tilde{s}$, the normalized radial distance $\tilde{r}$ and the normalized half-width $\tilde{r}_{1/2}$ as:
$\tilde{s}=\frac{s}{l},\quad\tilde{r}=\frac{r}{l},\quad\text{and}\quad\tilde{r}_{1/2}(s)=\frac{r_{1/2}(s)}{l}\;.$ (7)
@@ IV-A Experimental Setup
Figure 3: The DJI Matrice 300 drone hovers in proximity of the flow probe. The inset shows a close-up view of the Testo hotball probe. The flow probe measures the speed of the airflow.
Measurements of the flow around a drone require a large indoor space to avoid a recirculation of air while preventing external disturbances such as wind. A key enabler of this study is the access to a large industrial hall equipped with a 36-camera Vicon motion-capture system covering a tracking volume of $25\times 25\times 7\,\mathrm{m}$. This motion-capture system enables recording the position of the drone with millimeter accuracy.
The flow is measured with an omnidirectional hotball flowprobe (Testo 440). This type of commercial, off-the-shelf anemometer is indifferent to the flow direction and presents a sufficiently high measurement accuracy of ±$\pm 0.03\,\mathrm{m/s}\pm 5\%$ (up to 20 m/s) at a sampling rate of 1 s. This is favorable as we are interested in the induced mean magnitude and not the fluctuating velocity components of vortices in the turbulent flow. As shown in Fig. 3, the flow probe is mounted on a long pole at a height of 4.5 m, roughly corresponding to a normalized height of 5 tip-to-tip lengths of the drone.

### v2

# Robotics meets Fluid Dynamics: A Characterization of the Induced Airflow below a Quadrotor as a Turbulent Jet

@@ Abstract
The widespread adoption of quadrotors for diverse applications, from agriculture to public safety, necessitates an understanding of the aerodynamic disturbances they create. This paper introduces a computationally lightweight model for estimating the time-averaged magnitude of the induced flow below quadrotors in hover. Unlike related approaches that rely on expensive computational fluid dynamics (CFD) simulations or drone specific time-consuming empirical measurements, our method leverages classical theory from turbulent flows. By analyzing over 16 hours of flight data from drones of varying sizes within a large motion capture system, we show for the first time that the combined flow from all drone propellers is well-approximated by a turbulent jet after 2.5 drone-diameters below the vehicle. Using a novel normalization and scaling, we experimentally identify model parameters that describe a unified mean velocity field below differently sized quadrotors. The model, which requires only the drone’s mass, propeller size, and drone size for calculations, accurately describes the far-field airflow over a long-range in a very large volume which is impractical to simulate using CFD. Our model offers a practical tool for ensuring safer operations near humans, optimizing sensor placements and drone control in multi-agent scenarios. We demonstrate the latter by designing a controller that compensates for the downwash of another drone, leading to a four times lower altitude deviation when passing below.
@@ I Introduction
In recent years, quadrotors have gained popularity for a wide variety of tasks in academia [1, 2, 3] as well as in industry where companies develop drones for filming [4], mapping [5], inspection [6], and public safety [7]. Shared across these diverse applications is the need to understand and characterize the strong downwash generated by the vehicle’s propellers as this enables informed decisions on the allowed proximity of a quadrotor to an object or person, ultimately leading to safer autonomous drones.
@@
In mapping and inspection tasks this helps predicting when aerodynamic interactions with close-by structures such as bridges, powerlines [8], or ships [9] occur. When drones are operated for agricultural purposes, it is critical to know how far the aerodynamic disturbances caused by the drone extend for plant spraying and protection [10]. When deployed for filming and in public spaces, quadrotors are often operated in the vicinity of people where minimizing the presence of intrusive flows in the scene is important. In scenarios where multiple drones are operated together, a flow model can be used to improve planning such that individual vehicles dynamically avoid each other’s downwash and show an improved response to external disturbances [11]. Finally, a better understanding of the flow can be important for sensor and scientific instrumentation placement [12, 13, 14, 15, 16]. This paper presents a computationally lightweight approach to model the time-averaged magnitude of the induced flow below a quadrotor at hover.
Figure 1: Smoke visualization of the flow (left) and corresponding measured velocity field (right) of the Kolibri drone at hover. The individual propellers flows are separated close to the drone. After 2.5 motor-to-motor distances $l$ the individual flows merge into one turbulent jet (far-field) for which the velocity field and half-width (distance where the velocity is half the centerline velocity) can be calculated with our proposed method.
@@
Detailed computation of the induced flow around a drone is a challenging problem as it requires expensive and time-consuming computational fluid dynamics (CFD) simulations. To limit the computational demand, CFD simulations typically focus on millisecond time scales and simulate the airflow only close to the vehicle [17, 18, 19]. Such a simulation-based methodology is suitable to assess the drone flight performance, but not efficient in analyzing the far-field which extends over many meters away from the drone where one needs to rely on turbulence modeling [14, 16]. In addition, slow CFD simulations are not acceptable for a real-time planner or controller to dynamically coordinate a fleet of multiple quadrotors, such that the vehicles minimize interference with each other.
@@
Orthogonal to the simulation approach, fluid dynamics research has a century-old history of heavily relying on empirical measurement. However, experimentally characterizing each drone with thousands of measurements in the entire 3D space around it is a tedious process requiring a large, wind-free measurement domain. Additionally, a proper measurement technique needs to be selected. For example particle-image-velocimetry (PIV) [20] is well suited, but generally presents challenges in large-scale real-world deployments [21, 22].
@@
Our main contribution is a unified, computationally lightweight model to calculate the mean velocity of the induced flow below a drone at hover. Or, put differently, we answer the question: ‘How much wind does a drone generate when hovering or flying slowly in a near-hover state?’ The unified model is inspired by decades of well-established research on turbulent jets [23]. The advance in this work is enabled by joining methodologies from fluid dynamics and robotics research. Recording over 16 h of drone flight data with six differently sized drones ranging from 230 g to 6.3 kg we perform pointwise flow measurements in a very large motion capture system [24]. Through the use of appropriate normalization and the introduction of a characteristic drone length-scale the model is unified and can be applied to quadrotors of different mass and size.
Our methodology is computationally lightweight as it does not rely on CFD simulations. Instead, it uses available closed-form analytic flow solutions. All calculations only require knowledge of the vehicle’s mass, the vehicle’s dimensions, and the propeller size. We summarize our model in a pen-and-paper algorithm that calculates the downwash velocity field, making it easy to apply our findings to tasks in other domains, such as agriculture and public safety.
Additionally, to demonstrate the applicability of our simple method to real-world robotics applications, we integrate it into a controller that automatically compensates for downwash when passing below another drone, yielding a four times smaller altitude tracking error.
@@ II Related Work
For control and simulation tasks the primary goal of a model is to predict the motion of the vehicle given a certain actuation, that is, predict the forces and torques acting on its body. The most widespread models are rotor-based quadratic thrust and drag models [25, 26, 27] and more advanced models that are based on blade-element-momentum (BEM) theory which calculates the lift and drag at a propeller based on airfoil theory [28, 29, 30, 31, 32, 33, 34]. However, these approaches focus on estimating the aerodynamic forces and torques of the individual rotors but do not compute the combined total volumetric flow induced by the vehicle.
In contrast to the above models, computational fluid dynamics simulations do calculate the three-dimensional flow around the vehicle. Such approaches are typically used to determine and improve the efficiency of the design of a single propeller [35, 36], but have also been applied to simulate entire quadrotors [19, 18], and larger hexacopters [37, 14]. While such simulations achieve results that are highly accurate and manage to capture real-world effects [38, 39], they require large amounts of computation. Additionally, they do not easily generalize to other vehicles, in principle requiring a new analysis for each and every drone design.
In contrast to a purely simulation-based approach, a few works perform real-world experiments. For example, [12] performed flow probe measurements on a hexacopter that seem to find overall agreement with later CFD work [14]. In [40] the authors present experimental data to understand the influence of the quadrotor on atmospheric temperature and pressure measurements. In [41] the authors focus on the development of a Schlieren-photography method to qualitatively visualize the flow around drones. In [13] the authors performed outdoor smoke visualization, and in [42] the authors applied PIV to smoke visualizations to estimate the velocity field around a quadrotor model in ground effect.
Only a few other studies have applied PIV to drones. For example, in [43] the interaction between propellers under different configurations has been studied, and in [44] the near field flow of a quadrotor model in forward flight is analyzed using wind tunnel experiments. Most recently  [10] performed PIV in a different application and recovered the spreading rate of spray droplets in the drone’s downwash. Apart from [13, 40], all of these works performed measurements in tethered flight, presenting actual limitations in recovering the true free flight physics. To the authors’ best knowledge, no studies have looked into a comprehensive analysis of the flow below drones in hover while focusing on a scaling analysis, and on closed-form solutions from turbulent flows.
@@ III Theory
This section gives an overview of the relevant notation, overall drone thrust and propeller downwash in hovering flight, and introduces the concept of a turbulent jet flow. Finally, we introduce several normalizations enabling the analysis to generalize across a wide variety of drones.
Figure 2: Diagram for the drone and the flow coordinate system and geometry. The bodyframe $\mathcal{B}$ of a quadrotor is such that the x-axis faces forward, and the z-axis upwards in thrust direction. The flow coordinate system is oriented such that its longitudinal axis $\mathbf{\hat{s}_{\mathcal{F}}}$ is aligned with flow direction (i.e., points in negative $\mathbf{\hat{z}_{\mathcal{B}}}$ direction). The flow is described in cylindrical coordinates $[s~{}~{}r~{}~{}\theta]^{\top}$ by its longitudinal (axial), radial, and azimuthal velocity components $U$, $V$ and $W$, respectively. The vehicle’s propeller radius $r_{\text{prop}}$, diameter $d$ and motor-distance $l$ are defined as depicted.
@@ III-A Coordinate System and Notation
Figure 2 shows the quadrotor-centered bodyframe coordinate system $\mathcal{B}$, with vector $\mathbf{x}_{\mathcal{B}}=[x~{}~{}y~{}~{}z]^{\top}$. In the bodyframe the $\mathbf{\hat{x}_{\mathcal{B}}}$-axis points forward, the $\mathbf{\hat{y}_{\mathcal{B}}}$-axis points to the left, and the $\mathbf{\hat{z}_{\mathcal{B}}}$-axis points upwards. Since we only consider the quadrotor at hover, the $\mathbf{\hat{z}_{\mathcal{B}}}$-axis is opposing the gravity vector at all times and no rotation due to the roll and pitch of the vehicle needs to be considered. We also introduce a flow coordinate system $\mathcal{F}$ with $\mathbf{x}_{\mathcal{F}}=[s~{}~{}r~{}~{}\theta]^{\top}$ in cylindrical coordinates along the downward flow direction $\mathbf{\hat{s}_{\mathcal{F}}}$. The longitudinal coordinate $s$ is aligned with the downwash and the radial coordinate $r$ points outwards as shown in Fig. 2. In this coordinate system the velocity vector $\mathbf{U}$ with components $U,V,W$ is defined as shown in Fig. 2.
@@ III-B Propeller-Induced Flow
In a steady-state hover, the multicopter must produce enough thrust to support its weight. This means that, considering a quadrotor, the $N_{P}=4$ propellers must produce a thrust force that is equal to $mg$ where $m$ is the mass and $g$ is the gravitational acceleration.
@@
$U_{\rm{H}}=\sqrt{\frac{T_{\rm{H}}}{2\rho A_{\text{prop}}}}=\sqrt{\frac{\vphantom{T}mg}{2\rho\pi r_{\text{prop}}^{2}N_{P}}}\ ,$ (1)
@@
The air-density $\rho$ is computed according to the ideal gas law [45] to account for the local temperature $\vartheta$ and absolute pressure $p$ as
$\rho=pM/(R\vartheta)$ (2)
where $M=28.966\,\mathrm{g/mol}$ the molar mass of dry air and ${R=8.3144\,\mathrm{J/(K\cdot mol)}}$ is the universal gas constant.
@@ III-C Turbulent Jet Flow
For a turbulent jet, the mean longitudinal flow velocity $\langle U\rangle$ along the radial and flow-direction coordinates $r,s$ is given by the similarity profile [23]:
@@
Here $U_{\text{C}}$ describes the centerline velocity at $r=0$ as a function of the distance $s$ away from its exit nozzle, here below the rotor plane. For a jet with half-width $r_{1/2}$, the rescaled radial position $\xi$ is given by:
$\xi=r\,/\,r_{\raisebox{-1.5pt}{\scriptsize$1/2$}}(s)\ .$ (4)
The jet centerline velocity is known to scale inversely proportional to the distance from its exit. It is given by
@@
A key characteristic of the turbulent jet is that the spreading angle is commonly around 12 deg and is weakly dependent on the flow Reynolds number $Re_{\rm{d}}=\rho Ud/\mu$ that describes the relative importance between inertial and viscous forces [23]. This makes the model applicable to a wide range of drone flow and geometry, and independent of the air-viscosity $\mu$. We remark that we do not consider propeller swirl in the analysis as the rotors counter-rotate, and we assume that all other flow components are entrained (drawn/sucked into the flow from the side) downstream.
@@ III-D Normalization
To develop a generalized model of the airflow around different drones, the most important physical properties of those drones must be taken into account. We achieve this by introducing a velocity and a spatial normalization based on the theoretical considerations described previously.
@@
From (3) and (5) we observe that the overall scaling of the jet velocity is given by the jet-exit velocity $U_{\text{J}}$. The jet exit velocity describes how fast a jet exits a nozzle and captures how much momentum the flow carries. While a quadrotor does not have a nozzle, the induced velocity $U_{\rm{H}}$ of eq. (1) relates to the same physical quantity. Denoting a normalized quantity with $\sim$, the normalized velocity $\tilde{U}$ and normalized centerline velocity $\tilde{U}_{C}$ are given by:
$\langle\tilde{U}\rangle(\xi)=\langle U\rangle(\xi)/U_{\rm{H}}\quad\text{and}\quad\tilde{U}_{\text{C}}(s)=U_{\text{C}}(s)/U_{\rm{H}}\;.$ (7)
@@
Similar to the velocity normalization, a spatial normalization parameter is needed to be able to compare drones of different sizes. We propose the motor distance $l$ as the length-scale parameter as this distance is found to be the closest equivalent to a nozzle’s diameter in a jet flow. Using this, we define the normalized distance to the rotor plane $\tilde{s}$, the normalized radial distance $\tilde{r}$ and the normalized half-width $\tilde{r}_{1/2}$ as:
$\tilde{s}=s/l,\quad\tilde{r}=r/l,\quad\text{and}\quad\tilde{r}_{1/2}(s)=r_{1/2}(s)/l\;.$ (8)
@@ IV-A Experimental Setup
Figure 3: The DJI Matrice 300 drone hovers in proximity to the flow probe. The inset shows a close-up view of the Testo hot ball probe. The flow probe measures the speed of the airflow.
Measurements of the flow around a drone require a large indoor space to avoid air recirculation while preventing external disturbances such as wind. For this study a large industrial hall with a 36-camera Vicon motion-capture system ($25\times 25\times 7\,\mathrm{m}$ tracking volume) is used to record the drone’s position with millimeter accuracy.
The flow is measured with an omnidirectional hot ball flow probe (Testo 440) shown in Fig. 3. This type of commercial, off-the-shelf anemometer is indifferent to the flow direction and presents a sufficiently high measurement accuracy of ±$\pm 0.03\,\mathrm{m/s}\pm 5\%$ (up to 20 m/s) at a sampling rate of 1 s. This is favorable as we are interested in the induced mean magnitude and not the fluctuating velocity components of vortices in the turbulent flow.